OpenStudy (anonymous):

Solve the equation on the interval 0 ≤ θ < 2π. Solve the equation on the interval 0 ≤ θ < 2π. 2 cos2 θ - 1 = 0 it says that the answer its

OpenStudy (anonymous):

pi/4 3pi/4 5pi/4 and 7pi/4 Why?

OpenStudy (anonymous):

well, those aren't the answers..

OpenStudy (anonymous):

Can you explain the answers?

OpenStudy (anonymous):

2 cos2x -1 = 0 2 cos2x = 1 cos2x =1/2 => 2x = 60degrees or 150degrees => x = 30degrees or 75degrees...

OpenStudy (anonymous):

which one is the correct problem: \(\large 2cos^2\theta-1=0 \) OR \(\large 2cos(2\theta)-1=0 \) ????

OpenStudy (anonymous):

@dpaInc that's the matter..the first one is the equation..because it gives an answer from the answer choices

OpenStudy (anonymous):

This \[2\cos ^{2}\theta-1=0\]

OpenStudy (anonymous):

ok then.... to solve for theta, add 1 to both sides, divide by 2 to both sides, square root both sides.... then use the unit circle to determine what angles give you the necessary condition on the right side of the equation.